Un-polarized light passes through two polarizers;
Answer #1

a) Let initial intensity = Io

After one polarizer = Io/2

Final intensity = I = Io/2*cos^2 60 = 0.125*Io

b) I = 0.125*Io*cos^2 45 = 0.0625*Io